# Fully Kinetic Simulations of Proton-Beam-Driven Instabilities from Parker Solar Probe Observations

### v1

@@ Abstract
The expanding solar wind plasma ubiquitously exhibits anisotropic non-thermal particle velocity distributions. Typically, proton Velocity Distribution Functions (VDFs) show the presence of a core and a field-aligned beam. Novel observations made by Parker Solar Probe (PSP) in the innermost heliosphere have revealed new complex features in the proton VDFs, namely anisotropic beams that sometimes experience perpendicular diffusion. This phenomenon gives rise to VDFs that resemble a “hammerhead”. In this study, we use a 2.5D fully kinetic simulation to investigate the stability of proton VDFs with anisotropic beams observed by PSP. Our setup consists of a core and an anisotropic beam populations that drift with respect to each other. This configuration triggers a proton-beam instability from which nearly parallel fast magnetosonic modes develop. Our results demonstrate that before this instability reaches saturation, the waves resonantly interact with the beam protons, causing significant perpendicular heating at the expense of the parallel temperature. Furthermore, the proton perpendicular heating induces a hammerhead-like shape in the resulting VDF. Our results suggest that this mechanism probably contribute to producing the observed hammerhead distributions.
@@ 1 Introduction
For decades, spacecraft missions like Helios, Ulysses and Wind have conducted in situ observations, measuring velocity distribution functions (VDFs) of both ions and electrons in the solar wind (Marsch et al., 1982; Marsch, 2006). These observations revealed several features regarding the kinetic properties of the different particle species in the solar wind plasma. In this environment, not all plasma species have equal temperatures. Electrons are generally colder than protons in the fast solar wind but hotter than protons in the slow solar wind (Montgomery et al., 1968; Newbury et al., 1998). Proton populations can also present temperature anisotropy with respect to the direction of the local magnetic field (Bale et al., 2013). According to the double-adiabatic expansion model, in the absence of heat flux and collisions, protons in the inner heliosphere would show $T_{\perp,\mathrm{p}}<T_{\parallel,\mathrm{p}}$ (Chew et al., 1956). Here, $T_{\perp,\mathrm{p}}$ and $T_{\parallel,\mathrm{p}}$ denote the perpendicular and parallel temperatures of protons, respectively. However, observations reveal a predominance of cases where $T_{\perp,\mathrm{p}}>T_{\parallel,\mathrm{p}}$ in the fast solar wind within the inner heliosphere, underscoring the need for a mechanism driving substantial perpendicular heating (Richardson et al., 1995; Matteini et al., 2007). Proton VDFs are predominantly gyrotropic, often exhibiting skewness, resulting in a heat flux aligned along the magnetic field, directed away from the Sun (Marsch et al., 1982). They also often exhibit a magnetic field-aligned beam in the anti-sunward direction, a secondary proton population streaming at higher velocities compared to the core one (Asbridge et al., 1974; Alterman et al., 2018). It has been suggested that the process at the origin of the proton beam is the reflection of turbulent fluctuations at the MHD/kinetic spectral break: the greater the wavenumber of the spectral break, the faster the generated proton beam (Voitenko & Pierrard, 2015). These non-Maxwellian features are more pronounced in the fast solar wind compared to the slow solar wind, and their magnitude increases as the distance to the Sun decreases (Verscharen et al., 2019b). They represent symmetry-breaking properties capable of inducing plasma instabilities, leading to subsequent energy transfer processes restoring the equilibrium within the system (Verniero et al., 2020).
@@
The PSP mission (Fox et al., 2016) aims to study fundamental processes involved in energy transfer within the solar wind, such as acceleration and heating processes, to the solar surface. Observations from the Solar Wind Electron Alpha and Proton (SWEAP) instrument onboard PSP (Kasper et al., 2016) detected anisotropic beams in the proton VDFs, as reported by Klein et al. (2021). Moreover, extreme cases of anisotropic beam structures were also observed, specifically termed “hammerhead” (Verniero et al., 2022) due to their resemblance to the shape of a hammerhead. Consistently with the measurements of SWEAP, the FIELDS instrument (Bale et al., 2016) detected a strong wave activity from fluid MHD scales down to ion kinetic scales (Bowen et al., 2020; Verniero et al., 2020). The solar wind is a weakly collisional plasma (Alterman et al., 2018; Verscharen et al., 2019b), thus it is wave/particle interaction processes, rather than collisions, which act in the direction of reducing the non-Maxwellian character of VDFs (Kasper et al., 2002; Klein et al., 2018). The linear Alfvén dispersion relation at MHD scales connects with the fast-magnetosonic/whistler (FM/W) waves at kinetic scales (Belcher & Davis, 1971; Verscharen et al., 2019b). These modes are right-handed (RH) polarized and can be driven by temperature anisotropy with $T_{\perp}\ll T_{\parallel}$ in a high plasma beta, denoted as $\beta\doteq 8\pi nk_{\mathrm{B}}T/B^{2}\gg 1$ (Gary, 1993; Kennel & Engelmann, 1966). Here, $n$ represents the number density, $T$ signifies the temperature, $B$ denotes the magnetic field, and $k_{\mathrm{B}}$ stands for the Boltzmann constant. The observed relative drift speeds between core protons, secondary proton beams, and $\alpha$-particles (Marsch et al., 1982; Neugebauer et al., 1996; Steinberg et al., 1996; Podesta & Gary, 2011a, b) constitute another reservoir of free energy that can excite waves through wave-particle resonant interaction or cyclotron resonance (Bourouaine et al., 2013; Verscharen et al., 2013; Ofman et al., 2022, 2023). Damping mechanisms such as the cyclotron resonant damping (Goldstein et al., 1994; Leamon et al., 1998) dissipate energy at ion kinetic scales, where particles interact efficiently with electromagnetic waves (Leamon et al., 1998; Chandran et al., 2011). Through quasi-linear theory it is possible to detect signatures of cyclotron resonant damping in proton density phase-space (Hollweg & Isenberg, 2002; Cranmer, 2014; Bowen et al., 2022).
@@
The previously mentioned hammerhead structure (Verniero et al., 2022) is believed to originate from resonant interaction between parallel-propagating fluctuations and proton at kinetic scales, similar to those observed in (Klein et al., 2021). However, its precise origin has yet to be firmly established. Hybrid numerical simulation can efficiently study non-linear processes associated with unstable proton-beam VDFs (Ofman et al., 2022, 2023). However, in these codes, electrons are typically treated as a fluid, which leads to a loss of their fundamental kinetic nature. To gain insight into the mechanism behind the origin of the “hammerhead” (Verniero et al., 2022) from a first-principles approach and thus address the issue of perpendicular heating in solar wind plasma. To achieve this, we extend and complement the stability analysis of the proton-beam unstable VDF investigated by Klein et al. (2021) by exploring its nonlinear regime. In this study, we employ ECsim (Lapenta, 2017; Gonzalez-Herrero et al., 2019; Bacchini, 2023; Croonen, 2024), a semi-implicit, exactly energy-conserving PIC simulation tool, to investigate how the dynamics of both electrons and protons impact the development of the proton hammerhead. This paper is organized as follows: in Section 2, we present the setup of the PIC simulation in detail. Section 3 showcases the numerical modeling results alongside linear-theory analyses. Finally, Section 4 contains the discussion and conclusions.
@@ 2 Numerical setup
The non-Maxwellian protons, denoted by the subscript “p”, have constant number density $n_{\mathrm{p}}$, are divided into two sub-populations. The core, with subscript “c”, has number density $n_{\mathrm{c}}/n_{\mathrm{p}}=0.864$ and the beam, subscript “b”, has number density $n_{\mathrm{b}}/n_{\mathrm{p}}=0.136$. The core and the beam are both initialized with a bi-Maxwellian distribution, each characterized by a non-zero drift velocity along the $x$-axis (parallel direction):
@@
In equation (1) the subscript $i$ denotes the proton species, i.e. the core or the beam. The thermal velocities of the $i$-th proton species are expressed as $w_{i}\doteq\sqrt{k_{\mathrm{B}}T_{i}/m_{\mathrm{p}}}$, where $m_{\mathrm{p}}$ stands for the proton mass. Initially, these velocities are set to $w_{x,\mathrm{c}}/c_{\mathrm{A,p}}\approx 0.64$ and $w_{y,\mathrm{c}}/c_{\mathrm{A,p}}\approx 0.56$ for the core, while for the beam they are assigned as $w_{x,\mathrm{b}}/c_{\mathrm{A,p}}\approx 1.00$ and $w_{y,\mathrm{b}}/c_{\mathrm{A,p}}\approx 0.79$. The Alfvén speed for the protons defined as $c_{\mathrm{A,p}}\doteq B/\sqrt{4\pi n_{\mathrm{p}}m_{\mathrm{p}}}$ at the initialisation with a magnitude of $c_{\mathrm{A,p}}/c=0.00291$, where $c$ represents the speed of light in vacuum. Given that the plasma is described by a bi-Maxwellian distribution, there exists a symmetry in the perpendicular velocities, expressed as $w_{y,i}=w_{z,i}$ for all species $i$. The drift velocities of the core and beam are approximately $V_{x,\mathrm{c}}/c_{\mathrm{A,p}}\approx 0.18$ and $V_{x,\mathrm{b}}/c_{\mathrm{A,p}}\approx-1.34$, respectively. As a result, these two populations move towards each other. Electrons are initialized as a Maxwellian with thermal velocity $w_{\mathrm{e}}/c_{\mathrm{A,p}}\approx 7.12$ with unitary number density $n_{\mathrm{e}}$. The described electron-proton plasma satisfies the zero-net-current condition $n_{\mathrm{c}}V_{x,\mathrm{c}}+n_{\mathrm{b}}V_{x,\mathrm{b}}=0$ and the quasi-neutrality condition $n_{\mathrm{e}}=n_{\mathrm{p}}$.
@@
The proton VDF previously discussed is unstable and expected to produce fast magnetosonic modes, as demonstrated by the linear analysis conducted in (Klein et al., 2021). Based on linear theory calculations (for detailed analysis, see Section 3), the instability would manifest within a specific range of spatio-temporal scales, namely around those associated with the fastest growing mode. Consequently, we adjust the spatial and temporal parameters of our simulation accordingly. The fastest growing mode has a parallel wave number with a value of $k_{\parallel}/d_{\mathrm{p}}^{-1}\approx 0.5$ and a perpendicular one of $k_{\perp}/d_{\mathrm{p}}^{-1}\approx 0.03$. Here, $d_{\mathrm{p}}$ is the proton skin depth, defined as $d_{\mathrm{p}}\doteq c/\omega_{\mathrm{p}}$, and $\omega_{\mathrm{p}}\doteq\sqrt{4\pi e^{2}n_{\mathrm{p}}/m_{\mathrm{p}}}$ is the proton plasma frequency. This leads to a wavelength in the parallel direction of $\lambda_{x}/d_{\mathrm{p}}=2\pi/(k_{\parallel}/d_{\mathrm{p}}^{-1})\approx 13$ and in the perpendicular direction of $\lambda_{y}/d_{\mathrm{p}}=2\pi/(k_{\perp}/d_{\mathrm{p}}^{-1})\approx 209$. As a result, the domain length in each direction is selected to host several wavelength of this mode. This choice yields a rectangular box with dimensions $L_{x}/d_{\mathrm{p}}=64$ and $L_{y}/d_{\mathrm{p}}=256$. The spatial grid consists of $256\times 1024$ cells with a spatial increment of $\Delta x/d_{\mathrm{p}}=\Delta y/d_{\mathrm{p}}=0.25$. Similarly to the spatial domain, the estimation of the temporal scale is made by employing the growth rate of the most unstable mode $\gamma_{\mathrm{m}}/\omega_{\mathrm{p}}\approx 0.000121$ or $\gamma_{\mathrm{m}}/\Omega_{\mathrm{p}}\approx 0.042$, where $\Omega_{\mathrm{p}}\doteq eB_{0}/(m_{\mathrm{p}}c)$ is the proton cyclotron frequency. The ratio between the proton plasma and the cyclotron frequency is $\omega_{\mathrm{p}}/\Omega_{\mathrm{p}}\approx 344.04$. The instability evolves over a time scale $t/\omega_{\mathrm{p}}^{-1}=2\pi/(\gamma_{\mathrm{m}}/\omega_{\mathrm{p}})\approx 52000$. Therefore, the simulation is run for a total time $t_{\mathrm{tot}}/\omega_{\mathrm{p}}^{-1}\approx 150000$, to fully capture the development of the instability until the non-linear phase. The time step is set to $\Delta t/\omega_{\mathrm{p}}^{-1}=0.5$ to ensure accurate resolution of the proton gyro-period. We employed $4096$ particles per cell per species, initially distributed uniformly on the grid. We utilized a reduced proton-to-electron mass ratio, $m_{\mathrm{p}}/m_{\mathrm{e}}=183.6$, to narrow the gap between electron and proton dynamical scales, thereby saving computational time.

### v2

@@ Abstract
The expanding solar wind plasma ubiquitously exhibits anisotropic non-thermal particle velocity distributions. Typically, proton Velocity Distribution Functions (VDFs) show the presence of a core and a field-aligned beam. Novel observations made by Parker Solar Probe (PSP) in the innermost heliosphere have revealed new complex features in the proton VDFs, namely anisotropic beams that sometimes experience perpendicular diffusion. In this study, we use a 2.5D fully kinetic simulation to investigate the stability of proton VDFs with anisotropic beams observed by PSP. Our setup consists of a core and an anisotropic beam populations that drift with respect to each other. This configuration triggers a proton-beam instability from which nearly parallel fast magnetosonic modes develop. Our results demonstrate that before this instability reaches saturation, the waves resonantly interact with the beam protons, causing perpendicular heating at the expense of the parallel temperature.
@@ 1 Introduction
For decades, spacecraft missions like Helios, Ulysses and Wind have conducted in situ observations, measuring velocity distribution functions (VDFs) of both ions and electrons in the solar wind (Marsch et al., 1982; Marsch, 2006). These observations revealed several features regarding the kinetic properties of the different particle species in the solar wind plasma. In this environment, not all plasma species have equal temperatures. Electrons are generally colder than protons in the fast solar wind but hotter than protons in the slow solar wind (Montgomery et al., 1968; Newbury et al., 1998). Proton and electron populations can also present temperature anisotropy with respect to the direction of the local magnetic field (Bale et al., 2013). Temperature anisotropies in the solar wind can be generated by large-scale mechanisms such as solar wind expansion (Phillips & Gosling, 1990; Štverák et al., 2008; Opie et al., 2022), as well as by local wave-particle interactions (Tsurutani & Lakhina, 1997; Verscharen et al., 2013, 2019a). According to the double-adiabatic expansion model, in the absence of heat flux and collisions, protons in the inner heliosphere would show $T_{\perp,\mathrm{p}}<T_{\parallel,\mathrm{p}}$ (Chew et al., 1956). Here, $T_{\perp,\mathrm{p}}$ and $T_{\parallel,\mathrm{p}}$ denote the perpendicular and parallel temperatures of protons, respectively. However, observations reveal a predominance of cases where $T_{\perp,\mathrm{p}}>T_{\parallel,\mathrm{p}}$ in the fast solar wind within the inner heliosphere, underscoring the need for a mechanism driving substantial perpendicular heating (Richardson et al., 1995; Matteini et al., 2007). In Hellinger et al. 2011, 2013 and Matteini et al. 2013, the radial dependence of proton moments (e.g., density, bulk velocity, etc.) from Helios observations is fitted into the fluid equations for the parallel and perpendicular proton temperatures. These fits are used to calculate the heating (or cooling) rates in the parallel and perpendicular directions required to explain the observed radial trends in both the slow and the fast solar wind. That study finds that protons generally need to be heated perpendicularly between 0.3 and 1 AU, and that the necessary perpendicular proton heating rate is consistent with turbulent heating, as shown in Narita et al. 2022; Bowen et al. 2024. Proton VDFs are predominantly gyrotropic, often exhibiting skewness, resulting in a heat flux aligned along the magnetic field, directed away from the Sun (Marsch et al., 1982). They also often exhibit a magnetic field-aligned beam in the anti-sunward direction, a secondary proton population streaming at higher velocities compared to the core one (Asbridge et al., 1974; Alterman et al., 2018). It has been suggested that the process at the origin of the proton beam is the reflection of turbulent fluctuations at the MHD/kinetic spectral break: the greater the wavenumber of the spectral break, the faster the generated proton beam (Voitenko & Pierrard, 2015). These non-Maxwellian features are more pronounced in the fast solar wind compared to the slow solar wind, and their magnitude increases as the distance to the Sun decreases (Verscharen et al., 2019b). They represent symmetry-breaking properties capable of inducing plasma instabilities, leading to subsequent energy transfer processes restoring the equilibrium within the system (Verniero et al., 2020).
@@
The Parker Solar Probe (PSP) mission (Fox et al., 2016) aims to study fundamental processes involved in energy transfer within the solar wind and the solar Corona, such as acceleration and heating processes, to the solar surface. Observations from the Solar Wind Electron Alpha and Proton (SWEAP) instrument onboard PSP (Kasper et al., 2016) detected anisotropic beams in the proton VDFs, as reported by Klein et al. 2021. Moreover, extreme cases of anisotropic beam structures were also observed, specifically termed “hammerhead” (Verniero et al., 2022). Consistently with the measurements of SWEAP, the FIELDS instrument (Bale et al., 2016) detected a strong wave activity from Magnetohydrodynamics (MHD) scales down to ion kinetic scales (Bowen et al., 2020; Verniero et al., 2020). The solar wind is a weakly collisional plasma (Alterman et al., 2018; Verscharen et al., 2019b), thus it is wave/particle interaction processes, rather than collisions, which act in the direction of reducing the non-Maxwellian character of VDFs (Kasper et al., 2002; Klein et al., 2018). The linear Alfvén dispersion relation at MHD scales connects with the fast-magnetosonic/whistler (FM/W) waves at kinetic scales (Belcher & Davis, 1971; Verscharen et al., 2019b). Linear theory shows that, in the direction parallel to the background magnetic field ($\boldsymbol{k}\times\boldsymbol{B}_{0}=\boldsymbol{0}$), three different instabilities can arise in the presence of an ion core and a less dense ion beam population: the left-hand (LH) resonant, the non-resonant, and the right-hand (RH) resonant ion–ion instabilities (Gary, 1993). The observed relative drift speeds between core protons, secondary proton beams, and $\alpha$-particles (Marsch et al., 1982; Neugebauer et al., 1996; Steinberg et al., 1996; Podesta & Gary, 2011a, b) constitute another reservoir of free energy that can excite waves through wave-particle resonant interaction or cyclotron resonance (Bourouaine et al., 2013; Verscharen et al., 2013; Ofman et al., 2022, 2023). Indeed, proton beams, usually guided by the local magnetic field, have been observed in association with enhanced low-frequency electromagnetic fluctuations (Wicks et al., 2016; Klein et al., 2019; Martinović et al., 2021). These observations suggest that the fluctuations are produced by ion beam instabilities (Verscharen et al., 2019b; Zhu et al., 2023). Damping mechanisms such as cyclotron resonant damping (Goldstein et al., 1994; Leamon et al., 1998) dissipate energy at ion kinetic scales, where particles interact efficiently with electromagnetic waves (Leamon et al., 1998; Chandran et al., 2011). Through quasi-linear theory it is possible to detect signatures of cyclotron resonant damping in proton density phase-space (Hollweg & Isenberg, 2002; Cranmer, 2014; Bowen et al., 2022).
@@
The highly anisotropic beam structure analyzed by Verniero et al. 2022, referred to as the “hammerhead”, is believed to originate from resonant interaction between parallel-propagating fluctuations and protons at kinetic scales, similar to those observed in Klein et al. 2021. However, its precise origin has yet to be firmly established. Hybrid numerical simulations can efficiently study non-linear processes associated with unstable proton-beam VDFs (Ofman et al., 2022, 2023). However, in hybrid codes, electrons are treated as a fluid, which leads to a loss of their fundamental kinetic nature. We used a fully kinetic approach in which both protons and electrons are treated kinetically. This approach goes beyond the state of the art, where usually hybrid simulations are performed and provides a first verification of whether electron dynamics, in the specific physical scenario we consider, influences the issue of perpendicular proton heating in solar wind plasma. To achieve this, we extend and complement the stability analysis of the proton-beam unstable VDF investigated by Klein et al. 2021 by exploring its nonlinear regime. In this study, we employ ECsim (Lapenta, 2017; Gonzalez-Herrero et al., 2019; Bacchini, 2023; Croonen, 2024), a semi-implicit, exactly energy-conserving Particle-in-Cell (PIC) simulation tool, to investigate how the dynamics of both electrons and protons impact the development of the anisotropic proton beam. This paper is organized as follows: in Section 2, we present the setup of the PIC simulation in detail. Section 3 showcases the numerical modeling results alongside linear-theory analyses. Finally, Section 4 contains the discussion and conclusions.
@@ 2 Numerical setup
In our setup, the protons, denoted by the subscript “p”, have constant number density $n_{\mathrm{p}}$ and are divided into two sub-populations. The core, with subscript “c”, has number density $n_{\mathrm{c}}/n_{\mathrm{p}}=0.864$, while the beam, subscript “b”, has number density $n_{\mathrm{b}}/n_{\mathrm{p}}=0.136$. The core and the beam are both initialized with a bi-Maxwellian distribution, each characterized by a non-zero drift velocity along the $x$-axis (parallel direction):
@@
In equation (1), the subscript $i$ denotes the proton species, i.e. the core or the beam. The thermal velocities of the $i$-th proton species are expressed as $w_{i}\doteq\sqrt{k_{\mathrm{B}}T_{i}/m_{\mathrm{p}}}$, where $m_{\mathrm{p}}$ stands for the proton mass. Initially, these velocities are set to $w_{x,\mathrm{c}}/c_{\mathrm{A,p}}\approx 0.64$ and $w_{y,\mathrm{c}}/c_{\mathrm{A,p}}\approx 0.56$ for the core, while for the beam they are chosen as $w_{x,\mathrm{b}}/c_{\mathrm{A,p}}\approx 1.00$ and $w_{y,\mathrm{b}}/c_{\mathrm{A,p}}\approx 0.79$. The Alfvén speed for the protons is defined as $c_{\mathrm{A,p}}\doteq B_{0}/\sqrt{4\pi n_{\mathrm{p}}m_{\mathrm{p}}}$ at initialisation with a value of $c_{\mathrm{A,p}}/c=0.00291$, where $c$ represents the speed of light in vacuum. This results in the following temperature anisotropies: $T_{\perp,\mathrm{c}}/T_{\parallel,\mathrm{c}}=0.770$ and $T_{\perp,\mathrm{b}}/T_{\parallel,\mathrm{b}}=0.620$. Given that the plasma is described by a bi-Maxwellian distribution, there exists a symmetry in the perpendicular velocities, expressed as $w_{y,i}=w_{z,i}$ for all species $i$. The drift velocities of the core and beam are approximately $V_{x,\mathrm{c}}/c_{\mathrm{A,p}}\approx 0.18$ and $V_{x,\mathrm{b}}/c_{\mathrm{A,p}}\approx-1.16$, respectively. Electrons are initialized as a Maxwellian with thermal velocity $w_{\mathrm{e}}/c_{\mathrm{A,p}}\approx 7.12$ with unitary number density $n_{\mathrm{e}}$. The described electron-proton plasma satisfies the zero-net-current condition $n_{\mathrm{c}}V_{x,\mathrm{c}}+n_{\mathrm{b}}V_{x,\mathrm{b}}=0$ and the quasi-neutrality condition $n_{\mathrm{e}}\approx n_{\mathrm{p}}=n_{\mathrm{c}}+n_{\mathrm{b}}$.
@@
The proton VDF previously discussed is unstable and expected to produce fast magnetosonic modes, as demonstrated by the linear analysis conducted in Klein et al. 2021. Based on linear theory calculations (for a detailed analysis, see Section 3), the instability would manifest within a specific range of spatio-temporal scales, namely around those associated with the fastest growing mode. Consequently, we adjust the spatial and temporal parameters of our simulation accordingly. The fastest growing mode has a parallel wavenumber $k_{\parallel}/d_{\mathrm{p}}^{-1}\approx 0.501$ and a perpendicular one $k_{\perp}/d_{\mathrm{p}}^{-1}\approx 0.0245$. Here, $d_{\mathrm{p}}$ is the proton skin depth, defined as $d_{\mathrm{p}}\doteq c/\omega_{\mathrm{p}}$, and $\omega_{\mathrm{p}}\doteq\sqrt{4\pi e^{2}n_{\mathrm{p}}/m_{\mathrm{p}}}$ is the proton plasma frequency. This leads to a wavelength in the parallel direction $\lambda_{x}/d_{\mathrm{p}}=2\pi/(k_{\parallel}/d_{\mathrm{p}}^{-1})\approx 13$ and in the perpendicular direction $\lambda_{y}/d_{\mathrm{p}}=2\pi/(k_{\perp}/d_{\mathrm{p}}^{-1})\approx 209$. Therefore, the domain length in each direction is selected to fit several wavelengths of this mode. This choice yields a rectangular box with dimensions $L_{x}/d_{\mathrm{p}}=64$ and $L_{y}/d_{\mathrm{p}}=256$. The numerical grid consists of $256\times 1024$ cells with a spatial increment of $\Delta x/d_{\mathrm{p}}=\Delta y/d_{\mathrm{p}}=0.25$. Similarly to the spatial domain, the estimation of the temporal scale is made by employing the growth rate of the most unstable mode $\gamma_{\mathrm{m}}/\omega_{\mathrm{p}}\approx 0.000121$ or $\gamma_{\mathrm{m}}/\Omega_{\mathrm{p}}\approx 0.042$, where $\Omega_{\mathrm{p}}\doteq eB_{0}/(m_{\mathrm{p}}c)$ is the proton cyclotron frequency. The instability evolves over a time scale $t/\Omega_{\mathrm{p}}^{-1}=2\pi/(\gamma_{\mathrm{m}}/\Omega_{\mathrm{p}})\approx 151.15$. Therefore, the simulation is run for a total time $t_{\mathrm{tot}}/\omega_{\mathrm{p}}^{-1}\approx 436$, to fully capture the development of the instability until the non-linear phase. The time step is set to $\Delta t/\omega_{\mathrm{p}}^{-1}=0.00145$ to ensure accurate resolution of the proton gyro-period. We employed $4096$ particles per cell per species, initially distributed uniformly on the grid. We utilized a reduced proton-to-electron mass ratio, $m_{\mathrm{p}}/m_{\mathrm{e}}=183.6$, to narrow the gap between electron and proton dynamical scales, thereby saving computational time.